5) Find all absolute extrema of f(x)=x*-9x? +24x+10 on the interval (0,5).

Please be as detailed as possible thank you so much
Answer #1

f(x)= x² - 9x² +242 +10 [95] f(x) = 3x2 18x+24 f(x) = 6x-18 for critical point put f(x)=0 3x2_182424 50 3 (x2 - 6x+8)=0 x²_when x=2 then y=8-86 +48 +10=30 se (2130) when x=4 then = 64-144496 +10=26 jjc (4120) Now check extreme point at critical & ex=4 then f() = 26 x=5 then f() = 30 maximum at x = 5 so at x = 5 & 1=2 the given function has extreme point Absolute maximum